02. You are evaluating a peculiar real-estate venture that has the following cash flows: an outflow of IS210K at the end of the first year, a cash flow of IS98K in three years, a cash inflow of 1$101K two years after the preceding cash flow, a four-year annuity of cash flows with the first I$S70K cash flow occurring at t 9, and a perpetual series of $10K cash flows starting with the first cash flow 15 years from today. (a) Using a discount rate of 0.10, calculate the value of this real-estate opportunity. (b) According to these forecasted cash flows, what could you sell the venture for at t-142e same forecasted cash flows, what could you sel the venture for att 82 (d)Lastly, what could you sell the venture for at t 6?
Answer #1

For perpetuity cash flow that starts 15th year from today, Value at the end of the 14th year will be Cash flow / discount rate i.e. $10,000 / 10% = $100,000

So we can safely put +$100,000 in 14th year (in place for the perpetuity cash flow that starts in the 15th year)

Now, Various cash flows are shown below and Net present value is $280,218

t3 13 14 100,000 Year 91011 12 Cash Flows 210,000 98,000 98,000 98,000 101,000 101,000 Value 0,000 70,000 70,000 70,000 280,218 I-NPW 10%,C3P3)

b) We need to calculate the FV of the PV we calculated above after 14 years.

Fv = $280,218 * (1.1)^14 = $1,064,129.11

So we could sell the venture for $1,064,129.11 after 14 years.

c) At t=8, We could sell at $280,218 * 1.1^8 = $600,673.14

d) After 6 years, we could sell at $280,218 * 1.1^6 = $496,424.08